What is the negative solution of 2x^2+x=3?
Answer by stanbon(75887)

What is the negative solution of 2x^2+x=3?
-------
2x^2 + x - 3 = 0
---
Factor::
2x^2 +3x-2x-3 = 0
-----
x(2x+3)-(2x+3) = 0
-----
(2x+3)(x-1) = 0
-----
x = -3/2 or x = 1
------
Ans: x = -3/2
====================
Cheers,
Stan H.
-------------